# Cosmology from one galaxy in voids?

### v1

@@ 1 Introduction
Voids are sensitive probes to extract cosmological information (e.g. Park & Lee, 2007; Lavaux & Wandelt, 2010; Biswas et al., 2010; Lavaux & Wandelt, 2012; Pisani et al., 2015a; Hamaus et al., 2016; Pisani et al., 2019; Verza et al., 2019; Stopyra et al., 2021; Kreisch et al., 2021; Contarini et al., 2022a, b; Pelliciari et al., 2022; Schuster et al., 2023; Wang et al., 2023): for instance the void-galaxy cross-correlation function and the void size function already provide tight constraints on $\Omega_{\mathrm{m}}$ (Hamaus et al., 2020; Contarini et al., 2022b). Since voids are goldmines of cosmological information (Pisani et al., 2019; Moresco et al., 2022), it is interesting to ask whether galaxies inside voids bear a stronger constraining power from a cosmology perspective. A number of studies explored the properties of galaxies in voids (see e.g. Peebles, 2001; Jung et al., 2014; Habouzit et al., 2020; Panchal et al., 2020; Rosas-Guevara et al., 2022; Rodríguez Medrano et al., 2022; Rodríguez-Medrano et al., 2023b; Jaber et al., 2024). It has been shown, for example, that galaxies inside voids are expected to assemble stars slowly (Domínguez-Gómez et al., 2023). The observation of void galaxies is also an active field of research (see e.g. Grogin & Geller, 1999; Rojas et al., 2004; Hoyle et al., 2005; Croton et al., 2005; Kreckel et al., 2011, 2012; Moorman et al., 2015; Beygu et al., 2017; Douglass et al., 2018; Pustilnik et al., 2019; Florez et al., 2021; Pandey et al., 2021; Pustilnik et al., 2022; Kurapati et al., 2024), albeit some studies found contradictory results (for example related to the star formation rate and other properties of galaxies in voids). Due to their isolated nature, it is reasonable to expect that void galaxies—less influenced by external disturbances such as stripping and harassment—could have a stronger connection to cosmology and the physics of the early universe. Finally, differences have also been investigated in simulations, for example, more recently Rodríguez-Medrano et al. (2023a) and Curtis et al. (2024) suggested that void galaxies have significantly different properties compared to non-void galaxies (in the IllustrisTNG simulation), which may impact the link to cosmological information.
@@ 2.1 Simulations and Datasets
In this work, we used the dataset from the CAMELS project (Villaescusa-Navarro et al., 2021). Specifically, we use the galaxy catalogs from the LH set of the IllustrisTNG suite. The IllustrisTNG simulations are run with the AREPO code (Weinberger et al., 2020). This suite contains 1000 simulations of $(25h^{-1}Mpc)^{3}$ volume each and, while the values of $\Omega_{\rm b}=0.049$, $h=0.6711$, $n_{s}=0.9624$, $\sum m_{\nu}=0.0$ eV, $w=-1$ are fixed, the values of other cosmological and astrophysical parameters vary. In particular, the matter density parameter $\Omega_{\mathrm{m}}$, is varied within the range of 0.1 to 0.5. Notably, each simulation also varies four astrophysical parameters: two for supernovae ($A_{\rm SN1}$ and $A_{\rm SN2}$) and two for AGN ($A_{\rm AGN1}$ and $A_{\rm AGN2}$). Therefore, when inferring $\Omega_{\mathrm{m}}$, we implicitly marginalize over baryonic uncertainties. We followed the procedure described in Villaescusa-Navarro et al. (2022), and considered 17 galaxy properties (including $M_{\rm g}$, $M_{\rm BH}$, $M_{*}$, $M_{t}$, $V_{\rm{max}}$, $\sigma_{v}$, $Z_{\rm g}$, $Z_{*}$, SFR, $J$, $V$, $R_{*}$, $R_{\rm t}$, $R_{max}$ and the U, K, and g photometric bands for the CAMELS-IllustrisTNG galaxies, see more details in Villaescusa-Navarro et al., 2021). We consider a galaxy as a subhalo that contains no less than 20 particles.
@@ 2.2 The Void Finder
In this paper, we carefully need to select galaxies based on both their local environment and the large-scale environment. While the void center position (and any selection at a given distance from it) allows to account for the large-scale environment, the Voronoi cell volume guarantees a detailed galaxy selection that also accounts for their local environment. The distribution of the Voronoi cell volume for galaxies in our dataset is shown in Figure 1. Finally, void catalogs are subject to the presence of spurious voids (also known as Poisson voids) (Neyrinck, 2008; Pisani et al., 2015b; Cousinou et al., 2019). These are non-physical voids that arise due to a number of effects, including the minimum mass of the galaxies in the simulation. Generically spurious voids will present a shallower profile, with a higher density in the center and a smaller void size. Additional purity cuts need to be performed on the catalog to avoid bias from spurious voids.
A careful selection of galaxies is particularly relevant here, given the relatively small simulation box of $(25~{}h^{-1}{\rm Mpc})^{3}$ (compared to the size of cosmic voids, that usually span sizes in the range of a few to a hundred of ${}h^{-1}{\rm Mpc}$), and to account for the complex shape of voids. Details of how this galaxy selection is performed are provided in Section 2.4.
@@ 2.3 Predicting $\Omega_{\mathrm{m}}$ with neural networks using one galaxy
We follow Villaescusa-Navarro et al. (2022) and use fully connected layers with hyperparameters optimized by optuna (Akiba et al., 2019) to perform likelihood-free inference on $\Omega_{\mathrm{m}}$ using one galaxy. In order to fully utilize the dataset and train the networks to be adaptable to as many galaxies as possible, we still use all galaxies in the simulation to train our networks. Therefore, the training process is the same as Villaescusa-Navarro et al. (2022) and as a first test we have reproduced their results by using the public code available at https://github.com/franciscovillaescusa/Cosmo1gal. We use similar accuracy and precision metrics as in Echeverri-Rojas et al. (2023) to compare between void and non-void galaxies on predicting $\Omega_{\mathrm{m}}$. In particular, we rely on the rms error (RMSE), the mean relative error $\epsilon$ and the coefficient of determination $R^{2}$, defined as, respectively:
$\text{RMSE}=\sqrt{\langle(\theta-\mu)^{2}\rangle},$ (2)
$\epsilon=\left\langle\frac{\sigma}{\mu}\right\rangle,$ (3)
@@ 2.4 Void Galaxy Selection
After training the networks, we use the information from both voids and Voronoi cells inside voids to select void galaxies. As explained in Section 2.2, this combination allows us to jointly account for the large-scale structure and the local environment, providing a conservative selection of galaxies in isolated environments. Additionally, we consider purity cuts to avoid the presence of spurious voids. We note that our results are robust to variations in the presented cuts. We now provide details on our method to select void galaxies.
First, considering the void catalogs with no cuts, we begin with the set of galaxies belonging to VIDE voids. Since VIDE voids will include the overdense shells constituting the outer regions of voids, we perform a first selection of galaxies based on the distance from the void center, considering void galaxies as the ones within 75% of the void radius. The void radius is defined as $R=\left(\frac{3}{4\pi}V\right)^{1/3}$, where $V$ is the total volume of the Voronoi cells that the void contains. We note that, albeit this cut is based on a radius selection, the void center definition depends on the whole non-spherical shape of the void.
Second, to fully account for void shape for enhanced robustness, we rely on a criteria that is not based on radius, and utilize the Voronoi cell volume as an additional property to select void galaxies. Since larger Voronoi cell volumes correspond to more isolated galaxies, we consider different threshold values for Voronoi cell volumes to select void galaxies. The comparison of results with more or less stringent selection based on the Voronoi cell volumes is particularly instructive as it shows how results vary when considering more isolated galaxies. We show results with the various Voronoi cell volume thresholds in Section 3. In the remaining of the paper, we consider both the above criteria (cuts based on distance from the void center and on Voronoi cell volume) to separate void galaxies from non-void galaxies.
Finally, we add to the above selection additional purity cuts aimed at an enhanced robustness to the presence of spurious voids (Neyrinck, 2008; Pisani et al., 2015b; Cousinou et al., 2019). These purity cuts will aim to reduce the presence of galaxies belonging to spurious voids. Previous work showed that spurious voids are more likely to be voids of low size with a shallow density profile—that is a high density at their core (Cousinou et al., 2019). Therefore we exclude voids based on the void core density, defined as the density of the core particle in ZOBOV-normalized units (Neyrinck, 2008; Sutter et al., 2015), as well as on the void radius. These additional cuts allow to reduce the impact of spurious voids and provide a cleaner void catalog. We use each of these cuts to filter out $\sim 50\%$ of the voids from the catalogs. In Section 3 we show the results obtained when considering only a purity cut on void core density, and when considering both purity cuts on void core density and radius. We checked that our results are robust to variations in the aforementioned cuts.
@@ 3 Results
Figure 2: $\Omega_{\mathrm{m}}$ predictions using only one galaxy with fully connected layers. Blue dots and error bars represent the predicted posterior mean and standard deviation of $\Omega_{\mathrm{m}}$ from the networks. The orange line in the diagonal indicated the perfect match between predicted and true values. The left, middle and right columns show the results when using non-void galaxies, randomly selected galaxies and void galaxies respectively; therefore moving from left to right shows the change when considering the large-scale environment. For the right column, the top row shows the results after selecting void galaxies with a cut-off value of Voronoi cell volume at 100, while the bottom row shows a more restrictive Voronoi cell volume cut (150), corresponding to more isolated galaxies and accounting for the local environment. These results correspond to the void catalog considering the purity cut on core density. For better visual comparison, in the case of non-void and randomly selected galaxies we sub-sample the number of points to the same number of data points shown in the plot for void galaxies. The $R^{2}$ and RMSE values improve for void galaxies (from left to right), and further improve if we consider more isolated galaxies among the void galaxies (from top to bottom, right column).
@@
In this Section, we use the same trained machine learning model as in Villaescusa-Navarro et al. (2022), which takes galaxy properties as inputs and predicts the value of $\Omega_{\mathrm{m}}$ with fully connected layers. As described in Section 2.4, we have selected galaxies as void, non-void and randomly selected galaxies. We show the prediction of $\Omega_{\mathrm{m}}$ using these three types of galaxies. To investigate the impact of the galaxy environment on cosmological constraints, we consider galaxies at different level of isolation. This is done by considering two void galaxy selections determined by different cut-off values of the Voronoi cell volume, hence increasing how restrictive the selection is. In particular, we consider as void galaxies the galaxies that have Voronoi cell volumes larger than 100 and 150 (see Figure 1 for the overall range of volume values), with 150 being the most restrictive value, since a higher volume corresponds to more isolated galaxies, and show results for both cases.
Additionally, for the purity cuts at the level of the void catalog we show results after considering only the core density cut (Section 3.1), or both the core density cut and radius cut (Section 3.2). The results show the progressive improvement in constraints on $\Omega_{\mathrm{m}}$ with those cuts, indicating that a more restrictive void galaxy selection leads to better $\Omega_{\mathrm{m}}$ predictions. In Section 3.3, we show and discuss the average predictions of void, non-void and randomly selected galaxies in the same simulation.
@@ 3.1 Core Density Cut
In this Section, we only apply the core density cut to the void catalog. Figure 2 shows our results. Similar to Villaescusa-Navarro et al. (2022), the blue dots and error bars in Figures 2 and 3 represent the predicted posterior mean and standard deviation of $\Omega_{\mathrm{m}}$ from the networks. The orange diagonal line corresponds to the case in which the predicted values perfectly match the true values. The top left corners of each plot provide the values of accuracy and precision metrics. On the bottom right corners, we indicate the percentages of galaxies considered in the plots. Noticeably, for visualization purposes, in the case of non-void and randomly selected galaxies, we sub-sample the number of points to the same number of data points shown in the plot for void galaxies, allowing a better visual comparison. The metrics are, however, calculated over all the galaxies within each category. The left, middle, and right columns show the results of using non-void galaxies, randomly selected galaxies and void galaxies respectively. Therefore, moving from left to right shows the change when considering the large-scale environment. The first row in Figure 2 shows the results when using different types of galaxies to predict the values of $\Omega_{\mathrm{m}}$ with a less constrained (smaller) cell volume. From left to right in the row we can see that the result of using void galaxies is generally closer to the reference diagonal line, indicating that using void galaxies we can constrain $\Omega_{\mathrm{m}}$ better than when using non-void galaxies and randomly selected galaxies. The $R^{2}$ and RMSE values significantly improve when using void galaxies.
@@
This is consistent with the idea that, to guarantee an isolated environment, void galaxies need to be identified accounting for both the large-scale environment (accounted for when going from left to right in Figure 2) and the local environment (introduced when going from top to bottom in the right column of Figure 2). We note that in the bottom row, the data points in the figure might be visually less representative for the non-void and random galaxies cases, since for representation purposes we have sub-sampled the points in those cases to show the same number of points as for the void galaxies’ case. We note that in the bottom row the value of $\epsilon$ is not enhanced for void galaxies compared to randomly selected galaxies. This could be due to the smaller error bars and much lower number of data points in the case of more isolated galaxies (bottom row). The next Section shows results when considering both criteria to increase purity (core density cut and radius cut).

### v2

# Cosmology from one galaxy in a void?

@@ 1 Introduction
Voids are sensitive probes to extract cosmological information (e.g. Park & Lee, 2007; Lavaux & Wandelt, 2010; Biswas et al., 2010; Lavaux & Wandelt, 2012; Pisani et al., 2015a; Hamaus et al., 2016; Pisani et al., 2019; Verza et al., 2019; Stopyra et al., 2021; Kreisch et al., 2021; Contarini et al., 2022a, b; Pelliciari et al., 2022; Schuster et al., 2023; Wang et al., 2023): for instance the void-galaxy cross-correlation function and the void size function already provide tight constraints on $\Omega_{\mathrm{m}}$ (Hamaus et al., 2020; Contarini et al., 2022b). Since voids are goldmines of cosmological information (Pisani et al., 2019; Moresco et al., 2022), it is interesting to ask whether galaxies inside voids bear a stronger constraining power from a cosmology perspective.
Void galaxies are generally expected to be low-mass, blue, and with high star formation rates (Grogin & Geller (1999); Rojas et al. (2004, 2005); Hoyle et al. (2005); Park & Lee (2009); Kreckel et al. (2011); Florez et al. (2021). Some studies also suggest void galaxies may have lower metallicities (Pustilnik et al., 2003, 2005; Pustilnik & Tepliakova, 2011; Rosas-Guevara et al., 2022). However, whether void galaxies differ from comparable objects in denser regions remains controversial (see e.g. Kreckel et al. (2015); Penny et al. (2015); Habouzit et al. (2020)). The controversy can also be caused by a lack of a robust framework for studying void substructures and dynamics (Colberg et al., 2008). Still, due to their isolated nature, it is reasonable to expect that void galaxies—less influenced by external disturbances such as stripping and harassment—could have a stronger connection to cosmology and the physics of the early universe. In addition, differences have also been investigated specifically in the IllustrisTNG simulations recently: Rodríguez-Medrano et al. (2023) and Curtis et al. (2024) suggested that void galaxies have significantly different properties compared to non-void galaxies, which may impact the link to cosmological information.
@@ 2.1 Simulations and Datasets
In this work, we used the dataset from the Cosmology and Astrophysics with MachinE Learning Simulations (CAMELS) project, which includes more than 10,000 numerical simulations with different cosmological and astrophysical models (Villaescusa-Navarro et al., 2021). Specifically, we use the galaxy catalogs from the LH set of the IllustrisTNG suite. The IllustrisTNG simulations are run with the AREPO code (Weinberger et al., 2020). This suite contains 1000 simulations of $(25h^{-1}Mpc)^{3}$ volume each and, while the values of $\Omega_{\rm b}=0.049$, $h=0.6711$, $n_{s}=0.9624$, $\sum m_{\nu}=0.0$ eV, $w=-1$ are fixed, the values of other cosmological and astrophysical parameters are sampled from a latin hypercube (LH) with different initial random seeds. In particular, the matter density parameter $\Omega_{\mathrm{m}}$ is varied within the range of 0.1 to 0.5. Notably, each simulation also varies four astrophysical parameters: two for supernovae ($A_{\rm SN1}$ and $A_{\rm SN2}$) and two for AGN ($A_{\rm AGN1}$ and $A_{\rm AGN2}$). Therefore, when inferring $\Omega_{\mathrm{m}}$, we implicitly marginalize over baryonic uncertainties.
We follow the procedure described in Villaescusa-Navarro et al. (2022), and consider the following 17 galaxy properties:
  1. 1.
    $M_{\rm g}$: The mass of gas in the galaxy, including contributions from the circumgalactic medium.
  2. 2.
    $M_{\rm BH}$: The mass of the galaxy’s black hole.
  3. 3.
    $M_{*}$: The stellar mass of the galaxy.
  4. 4.
    $M_{\rm t}$: The total mass of the subhalo hosting the galaxy, comprising dark matter, gas, stars, and black holes.
  5. 5.
    $V_{\rm max}$: The maximum circular velocity of the galaxy’s subhalo, defined as $V_{\rm max}=\max(\sqrt{GM(<R)/R}$).
  6. 6.
    $\sigma_{v}$: The velocity dispersion of all particles within the galaxy’s subhalo.
  7. 7.
    $Z_{\rm g}$: The mass-weighted metallicity of the galaxy’s gas.
  8. 8.
    $Z_{*}$: The mass-weighted metallicity of the galaxy’s stars.
  9. 9.
    ${\rm SFR}$: The star formation rate of the galaxy.
  10. 10.
    $J$: The magnitude of the spin vector of the galaxy’s subhalo.
  11. 11.
    $V$: The magnitude of the peculiar velocity of the galaxy’s subhalo.
  12. 12.
    $R_{*}$: The radius that contains half of the galaxy’s stellar mass.
  13. 13.
    $R_{\rm t}$: The radius that contains half of the total mass of the galaxy’s subhalo.
  14. 14.
    $R_{\rm max}$: The radius at which $\sqrt{GM(<R_{\rm max})/R_{\rm max}}=V_{\rm max}$.
  15. 15.
    ${\rm U}$: The galaxy’s magnitude in the U band.
  16. 16.
    ${\rm K}$: The galaxy’s magnitude in the K band.
  17. 17.
    ${\rm g}$: The galaxy’s magnitude in the g band.
In addition, we also consider a galaxy as a subhalo that contains no less than 20 particles as stated in Villaescusa-Navarro et al. (2022). These galaxies have masses ranging from $\sim 10^{8}\mathrm{M}_{\odot}$ to $\sim 10^{14}\mathrm{M}_{\odot}$ with $\sim 80\%$ of galaxies under $\sim 10^{11}\mathrm{M}_{\odot}$ and $\sim 25\%$ of galaxies under $\sim 10^{10}\mathrm{M}_{\odot}$.
We follow Villaescusa-Navarro et al. (2022) and use fully connected layers with hyperparameters optimized by optuna (Akiba et al., 2019) to perform likelihood-free inference on $\Omega_{\mathrm{m}}$ using one galaxy. In order to fully utilize the dataset and train the networks to be adaptable to as many galaxies as possible, we still use all galaxies in the simulation to train our networks. Therefore, the training process is the same as Villaescusa-Navarro et al. (2022) and as a first test we have reproduced their results by using the public code available at https://github.com/franciscovillaescusa/Cosmo1gal. We use similar accuracy and precision metrics as in Echeverri-Rojas et al. (2023) to compare between void and non-void galaxies on predicting $\Omega_{\mathrm{m}}$. In particular, we rely on the root mean square error (RMSE), the mean relative error $\epsilon$ and the coefficient of determination $R^{2}$, defined as
$\text{RMSE}=\sqrt{\langle(\theta-\mu)^{2}\rangle},$ (2)
$\epsilon=\left\langle\frac{\sigma}{\mu}\right\rangle,$ (3)
@@ 2.3 The Void Finder
In this paper, we carefully need to select galaxies based on both their local environment and the large-scale environment. While the void center position (and any selection at a given distance from it) allows to account for the large-scale environment, the Voronoi cell volume guarantees a detailed galaxy selection that also accounts for their local environment. The distribution of the Voronoi cell volume for galaxies in our dataset is shown in Figure 1.
A careful selection of void galaxies is particularly relevant here, given the relatively small simulation box of $(25~{}h^{-1}{\rm Mpc})^{3}$ (compared to the size of cosmic voids, that usually span sizes in the range of a few to a hundred of ${}h^{-1}{\rm Mpc}$), and to account for the complex shape of voids. Details of how this selection is performed are provided in the following section.
@@ 2.4 Void Galaxy Selection
We jointly account for the large-scale structure and the local environment and perform a conservative selection of void galaxies in isolated environments. Additionally, we consider purity cuts to avoid the presence of spurious voids.
First, considering the void catalogs with no cuts, we begin with the set of galaxies belonging to VIDE voids. Since VIDE voids will include the overdense shells constituting the outer regions of voids, we perform a first selection of galaxies based on the distance from the void center, considering void galaxies as the ones within 75% of the void radius. The void radius is defined as $R=\left(\frac{3}{4\pi}V\right)^{1/3}$, where $V$ is the total volume of the Voronoi cells that the void contains. We note that, although this cut is based on a radius selection, the void center definition depends on the whole non-spherical shape of the void.
Second, to fully account for void shape for enhanced robustness, we rely on a criteria that is not based on radius, and utilize the Voronoi cell volume as an additional property to select void galaxies. Since larger Voronoi cell volumes correspond to more isolated galaxies, we consider different threshold values for Voronoi cell volumes to select void galaxies. The comparison of results with more or less stringent selection based on the Voronoi cell volumes is particularly instructive as it shows how results vary when considering more isolated galaxies. We show results with the various Voronoi cell volume thresholds in Section 3. In the remainder of the paper, we consider both the above criteria (cuts based on distance from the void center and on Voronoi cell volume) to separate void galaxies from non-void galaxies.
Finally, we add to the above selection additional purity cuts aimed at an enhanced robustness to the presence of spurious voids (Neyrinck, 2008; Pisani et al., 2015b; Cousinou et al., 2019). These purity cuts aim to reduce the presence of galaxies belonging to spurious voids. Previous work showed that spurious voids are more likely to be voids of low size with a shallow density profile—that is a high density at their core (Cousinou et al., 2019). Therefore we exclude voids based on the void core density, defined as the density of the core particle in ZOBOV-normalized units (Neyrinck, 2008; Sutter et al., 2015), as well as on the void radius. These additional cuts allow us to reduce the impact of spurious voids and provide a cleaner void catalog. We use each of these cuts to filter out $\sim 50\%$ of the voids from the catalogs. In Section 3 we show the results obtained when considering only a purity cut on void core density, and when considering both purity cuts on void core density and radius. We checked that our results are robust to variations in the aforementioned cuts.
@@ 3 Results
Figure 2: $\Omega_{\mathrm{m}}$ predictions using only one galaxy with fully connected layers. Blue dots and error bars represent the predicted posterior mean and standard deviation of $\Omega_{\mathrm{m}}$ from the networks. The orange line in the diagonal indicated the perfect match between predicted and true values. The left, middle and right columns show the results when using non-void galaxies, randomly selected galaxies and void galaxies respectively; therefore moving from left to right shows the change when considering the large-scale environment. For the right column, the top row shows the results after selecting void galaxies with a cut-off value of Voronoi cell volume at 100 $(\mathrm{Mpc}/h)^{3}$, while the bottom row shows a more restrictive Voronoi cell volume cut at 150 $(\mathrm{Mpc}/h)^{3}$, corresponding to more isolated galaxies and accounting for the local environment. These results correspond to the void catalog considering the purity cut on core density. For better visual comparison, in the case of non-void and randomly selected galaxies we sub-sample the number of points to the same number of data points shown in the plot for void galaxies. The $R^{2}$ and RMSE values improve for void galaxies (from left to right), and further improve if we consider more isolated galaxies among the void galaxies (from top to bottom, right column).
@@
In this Section, we use the same trained machine learning model as in Villaescusa-Navarro et al. (2022), which takes galaxy properties as inputs and predicts the value of $\Omega_{\mathrm{m}}$ with fully connected layers. As described in Section 2.4, we have classified galaxies into three groups: void, non-void, and random galaxies. Random galaxies refer to galaxies chosen without any specific criteria, irrespective of the environment. We show the prediction of $\Omega_{\mathrm{m}}$ using these three types of galaxies. To investigate the impact of the galaxy environment on cosmological constraints, we consider galaxies at different levels of isolation. This is done by considering two void galaxy selections determined by different cut-off values of the Voronoi cell volume, hence increasing how restrictive the selection is. In particular, we consider as void galaxies the galaxies that have Voronoi cell volumes larger than 100 $(\mathrm{Mpc}/h)^{3}$ and 150 $(\mathrm{Mpc}/h)^{3}$ (see Figure 1 for the overall range of volume values), with 150 $(\mathrm{Mpc}/h)^{3}$ being the most restrictive value, since a higher volume corresponds to more isolated galaxies, and show results for both cases.
Additionally, for the purity cuts at the level of the void catalog we show results after considering only the core density cut (Section 3.1), or both the core density cut and radius cut (Section 3.2). The results show the progressive improvement in constraints on $\Omega_{\mathrm{m}}$ with those cuts, indicating that a more restrictive void galaxy selection leads to better $\Omega_{\mathrm{m}}$ predictions. In Section 3.3, we show and discuss the average predictions of void, non-void and randomly selected galaxies to investigate the constraining capabilities of the different kinds of galaxies separately.
@@ 3.1 Core Density Cut
In this Section, we only apply the core density cut to the void catalog. Figure 2 shows our results. Similar to Villaescusa-Navarro et al. (2022), the blue dots and error bars in Figures 2 and 3 represent the predicted posterior mean and standard deviation of $\Omega_{\mathrm{m}}$ from the networks. The orange diagonal line corresponds to the case in which the predicted values perfectly match the true values. The top left corners of each plot provide the values of accuracy and precision metrics. On the bottom right corners, we indicate the percentages of galaxies considered in the plots. Noticeably, for visualization purposes, in the case of non-void and randomly selected galaxies, we sub-sample the number of points to the same number of data points shown in the plot for void galaxies, allowing a better visual comparison. The metrics are, however, calculated over all the galaxies within each category. The left, middle, and right columns show the results of using non-void galaxies, randomly selected galaxies and void galaxies respectively. Therefore, moving from left to right shows the change when considering the large-scale environment. The first row in Figure 2 shows the results when using different types of galaxies to predict the values of $\Omega_{\mathrm{m}}$ with a less constrained (smaller) cell volume. From left to right in the row we can see that the result of using void galaxies is generally closer to the reference diagonal line, indicating that using void galaxies we can constrain $\Omega_{\mathrm{m}}$ better than when using non-void galaxies and randomly selected galaxies. The $R^{2}$ and RMSE values improve $\sim 5\%$ and $\sim 14\%$ respectively when using void galaxies.
@@
This is consistent with the idea that, to guarantee an isolated environment, void galaxies need to be identified accounting for both the large-scale environment (accounted for when going from left to right in Figure 2 and the local environment (introduced when going from) (accounted for when going from left to right in Figure 2) and the local environment (introduced when going from top to bottom in the right column of Figure 2). We note that in the bottom row, the data points in the figure might be visually less representative for the non-void and random galaxies cases, since for representation purposes we have sub-sampled the points in those cases to show the same number of points as for the void galaxies’ case. We also notice that the value of $\epsilon$ is not enhanced for void galaxies compared to randomly selected galaxies. This could be due to the smaller error bars and much lower number of data points in the case of more isolated galaxies (bottom row). The next Section shows results when considering both criteria to increase purity (core density cut and radius cut).